# An Interdisciplinary and Cross-Task Review on Missing Data Imputation

### v1

@@ 1 Introduction
Missing data is a prevalent challenge across numerous fields, including social science, e-commerce, healthcare, bioinformatics, and communications. For instance, in social science surveys, respondents may refuse to answer certain questions, leading to non-responses (Little and Rubin, 1989). In bioinformatics, gene expression matrices often contain significant missing values, or ‘dropouts’, which occur when expressed transcripts are not fully detected and are consequently recorded as zeros (Troyanskaya et al., 2001; Huang et al., 2018). Similarly, in sensor networks, node failures can result in incomplete data collection (Faizin et al., 2019). In e-commerce, the user-item interaction matrix is typically highly sparse, and predicting its missing values forms the basis of recommendation systems (Koren et al., 2009). Even in computer vision, tasks such as watermark and occlusion removal can be framed as missing data imputation problems. Beyond imputation, two other primary approaches for handling missing data are deletion and ignorance. The deletion method involves discarding any data instances or features that contain missing values, proceeding with analysis only on the remaining complete dataset. A common implementation is to remove rows or columns with missing entries from a data matrix. A significant drawback of this approach is that it can drastically reduce the dataset size, potentially compromising the reliability of subsequent analysis. Moreover, it fails to leverage the partial information available in the incomplete observations. The ignorance approach, conversely, overlooks the presence of missing values during analysis, often by replacing them with a simple placeholder like zero. This method presupposes that the analytical or learning algorithms are robust to such missingness. Intuitively, this strategy is only viable when the proportion of missing data is sufficiently small.
@@
Research on handling missing data dates back to the 1930s (Wilks, 1932; Yates, 1933). For instance, Wilks (1932) investigated the estimation of means, variances, and covariances for a bivariate normal population from incomplete samples, comparing two methods: maximum likelihood estimation and an approach based on systems of independent estimates. This foundational work was later expanded by (Lord, 1955) and (Anderson, 1957), who addressed the more general problem of estimating parameters for multivariate normal populations with incomplete data. A key literature review on this topic in multivariate statistics was subsequently compiled by (Afifi and Elashoff, 1966). A significant theoretical advance was made by (Rubin, 1976), who established that directly ignoring the missing data process is inappropriate for likelihood-based inference unless the data are Missing at Random. Around the same time, Dempster et al. (1977) introduced the Expectation-Maximization (EM) algorithm for obtaining maximum likelihood estimates from incomplete data, which has since become one of the most influential methods in the field. Parallel to these methodological developments, researchers such as Kalton and Kasprzyk (1982), Little (1988), and Little and Rubin (1989) focused on missing data problems in the social sciences. Among them, Little and Rubin (1989) provided a detailed analysis of different missingness patterns, including general and special cases like univariate and monotone missingness. From the perspective of psychological research, Roth (1994) later reviewed classical methods for missing data imputation.
@@
Since real-world data are often represented in matrix form, several scholars have proposed performing missing data imputation by leveraging the underlying structure of the matrix, a problem known as matrix completion (Candès and Recht, 2009; Mazumder et al., 2010). A particularly important and useful structure is low-rankness, which arises from the fact that real data matrices often lie in a low-dimensional subspace (Udell and Townsend, 2019). This low-rank prior enables the effective recovery of missing values. While numerous approaches to matrix completion exist, some researchers have extended the low-rank assumption to more general structures, such as unions of subspaces or manifolds, which can lead to high-rank or even full-rank matrices (Eriksson et al., 2011; Fan and Chow, 2018; Ongie et al., 2017; Fan et al., 2020a). For instance, Fan et al. (2020a) assumed that the data are drawn from a union of polynomials. Matrix completion methods have found applications in data preprocessing, image inpainting, collaborative filtering (and recommendation systems more broadly) (Koren et al., 2009), and link prediction (Menon and Elkan, 2011).
@@
Higher-order tensors, as natural extensions of matrices, are also ubiquitous in science and engineering. For example, a color image with three channels can be represented as a third-order tensor, and fMRI scans across multiple time points can form a fourth-order tensor. To impute missing values in such data, tensor completion has been extensively studied as an extension of matrix completion (Gandy et al., 2011; Acar et al., 2011; Liu et al., 2012; Qin et al., 2022). Tensor completion has applications in recommendation systems, image and video inpainting, and knowledge graph completion, among others. Compared to matrix completion, tensor completion typically involves higher time and space complexity, and the landscape of tensor decomposition models is more diverse. However, both theoretical and empirical studies have shown that tensor completion often significantly outperforms matrix completion applied to matricized tensors in terms of recovery accuracy.
@@
Recently, deep learning (LeCun et al., 2015; Goodfellow et al., 2016), a prominent branch of machine learning, has been increasingly applied to missing data imputation. Deep learning typically refers to deep artificial neural networks, which are capable of approximating highly complex functions, as established by the universal approximation theorems (Pinkus, 1999; Sonoda and Murata, 2017; Lu et al., 2017). This capacity enables deep learning to capture the intricate structures inherent in real-world data, leading to impressive performance in imputation tasks. Perhaps the most well-known deep learning model for this purpose is the autoencoder (Hinton and Salakhutdinov, 2006). In this architecture, the input typically consists of incomplete data, and the output is the reconstructed, complete data (Fan and Chow, 2017a; Gondara and Wang, 2018; Xie et al., 2020). Beyond autoencoders, the widely recognized generative adversarial network (GAN) (Goodfellow et al., 2014) has also been adapted for imputation (Yoon et al., 2018; Li et al., 2019; Yoon and Sull, 2020). For instance, GAIN, introduced by Yoon et al. (2018), incorporates additional ‘hint’ information into the discriminator to enable its operation on incomplete data. These hints guide the generator to produce imputations that adhere to the true underlying data distribution. Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Yang et al., 2023) have also been leveraged for missing data imputation (Tashiro et al., 2021). More recently, the potential of Large Language Models (LLMs) for missing data imputation has been investigated. More details will be provided in Section 5.7.
@@ 2 Related Work
The period from 2000 to 2010 yielded several influential surveys on missing data. Specifically, Pigott (2001) summarized common methods, including complete-case analysis, available-case analysis, single imputation, and model-based methods for multivariate normal data, as well as maximum likelihood via the EM algorithm. Schafer and Graham (2002) provided a detailed discussion of the Missing at Random assumption and evaluated the effectiveness of maximum likelihood and Bayesian multiple imputation. During this time, Little and Rubin (2002) also published a seminal book on the statistical analysis of missing data, the most recent edition of which was released in 2019. Meanwhile, Donders et al. (2006) offered a concise review that compared single and multiple imputation through numerical experiments. Graham (2009) presented a practical overview of the missing data literature, covering theoretical foundations and methods like normal-model multiple imputation and maximum likelihood, while also addressing practical issues such as using auxiliary variables to enhance power and reduce bias. The subsequent decade saw a continuation of this review effort. In 2010, Enders C.K. authored an applied book on handling missing data, with a latest edition published in 2022 (Enders, 2022). That same year, Baraldi and Enders (2010) explored the theoretical foundations of missing data analysis, summarized conventional techniques, and provided accessible explanations of maximum likelihood and multiple imputation. After 2010, there are also a few review papers on missing data (Nakai and Ke, 2011; Eekhout et al., 2012; Silva and Zárate, 2014; Shen et al., 2015; Pratama et al., 2016; Gabrio et al., 2017; Wu et al., 2019a). For instance, Pratama et al. (2016) briefly reviewed a few methods for handling missing data in time series.
A number of recent reviews have explored imputation methods across various fields. Zhang and Thorburn (2022) surveyed methods for environmental monitoring, while Sethia et al. (2023) provided a brief overview that categorizes techniques into simple and multiple imputation. In a 2024 review of 46 studies on electronic health records, Ren et al. (2024a) concluded that while machine learning methods show significant promise, no single approach offers a universally generalizable solution. More recently, Chourib (2025) presented a short review that classified strategies into three principal categories: preprocessing techniques, graph-based imputation, and algorithms inherently tolerant to missing values, accompanied by a numerical evaluation of five methods. It is worth noting that numerous other publications are titled as reviews or surveys on missing data, but most of them have a very narrow focus (e.g., (Moorthy et al., 2014; Faizin et al., 2019; Wu et al., 2019a; Zhang et al., 2025b)) and will not be discussed here.
@@
(Roth, 1994) psychology
(Pigott, 2001) statistics
(Troyanskaya et al., 2001) bioinformatics
(Schafer and Graham, 2002) statistics
(Little and Rubin, 2002) statistics
(Graham, 2009) psychology
(Porter and Ecklund, 2012) social science
(Cheema, 2014) education
(Silva and Zárate, 2014) data analytics
(Shen et al., 2015) remote sensing
(Pratama et al., 2016) time series
(Du et al., 2020) monitoring systems
(Emmanuel et al., 2021) machine learning
(Thomas and Rajabi, 2021) machine learning
(Joel et al., 2022) machine learning
(Adhikari et al., 2022) internet of things
(Nijman et al., 2022) medical
(Liu et al., 2023a) healthcare & machine learning
(Sun et al., 2023) machine learning
(Ren et al., 2024a) healthcare & machine learning
(Alwateer et al., 2024) machine learning
(Zhang et al., 2024c) transportation
(Little, 2024) psychology
(Le et al., 2025) healthcare
(Benhamza et al., 2025) healthcare
(Chourib, 2025) healthcare
@@ 3.1 Mechanisms of Missingness
When data are classified as missing at random (MAR), the occurrence of missing data is systematically linked to the observed data, but not to the unobserved data (Rubin, 1976; Heitjan and Basu, 1996; Little and Rubin, 2019). More formally, following the previous definition of $\mathbf{X}$ and $\mathbf{M}$, we split $\mathbf{X}$ into two parts, i.e., $\mathbf{X}=(\mathbf{X}_{\mathrm{obs}},\mathbf{X}_{\mathrm{mis}})$, where $\mathbf{X}_{\mathrm{obs}}$ and $\mathbf{X}_{\mathrm{mis}}$ denote the observed values and missing values respectively. The data are MAR if $P(\mathbf{M}\mid\mathbf{X})=P(\mathbf{M}\mid\mathbf{X}_{\mathrm{obs}})$. For instance, in a survey, if men are more likely to withhold their weight than women, but this non-disclosure is not connected to the actual weights, then such data would be considered MAR. In other words, the likelihood of survey completion is associated with their gender (which is completely observed) but not their weight. MAR is a weaker assumption than MCAR and is more common in real-world data. However, handling MAR data is more complex and usually requires more sophisticated statistical techniques to avoid bias, such as multiple imputation or maximum likelihood estimation.
@@
Missing not at random (MNAR), also referred to as non-ignorable nonresponse, represents a type of data that is neither MAR nor MCAR. In instances where data is MNAR, the absence of data is systematically connected to the unseen data, meaning that the missingness is linked to events or factors that the researcher has not measured. To elaborate with the previous example, a weight registry might face MNAR data if participants who are overweight are more inclined to decline completing a weight survey. Another instance could be a survey including sensitive queries like personal income. If individuals with higher earnings are more prone to withhold their income information (resulting in missing data), the missingness of the income data correlates with the actual missing income values. This situation exemplifies MNAR. More details and subtypes of MNAR can be found in (Little and Rubin, 2019; Pereira et al., 2024). Addressing MNAR is a challenging task and may yield skewed results if not appropriately managed.
@@ 3.2 Single Imputation and Multiple Imputation
In single imputation, each missing value is replaced with a single estimated value. The methods used to estimate the missing value can vary, including mean imputation, median imputation, regression imputation, and many others that will be discussed in the subsequent sections of this paper. While single imputation is simple and easy to implement, it has significant limitations. It doesn’t reflect the uncertainty about the imputations (since each missing value is replaced by a single value), which can lead to underestimated standard errors and overly confident statistical inferences (Rubin, 1987).
Multiple imputation addresses the limitations of single imputation by replacing each missing value with a set of plausible values, creating multiple complete datasets. These datasets are then analyzed separately, and the results are combined to produce estimates and confidence intervals that incorporate missing-data uncertainty. This process reflects the uncertainty about the right model to impute missing data and the randomness inherent in the data itself. As a result, multiple imputation provides more accurate and statistically valid results than single imputation (Rubin, 2018, 1996; Murray, 2018). The book (Rubin, 2018) provided a very detailed introduction, application, and discussion of multiple imputation. Compared to single imputation, the major limitation of multiple imputation is the higher computation cost, especially when advanced imputation algorithms are used.
@@ 3.3 Goals of Missing Data Imputation
Missing data imputation not only plays an important role in data preprocessing, but also is the final goal of many real scenarios such as recommendation systems. Moreover, missing data imputation can be used to improve efficiency and reduce the cost of data acquisition and analysis.
@@
In many data analysis tasks—such as classification, regression, clustering, novelty detection, and causal inference—algorithms typically require complete data. The performance of these algorithms is often directly influenced by the quality of the imputed values. For example, a support vector machine trained on data where missing values are zero-filled might achieve 60% classification accuracy on a test set; this accuracy can improve to 80% when a reliable imputation algorithm is used. The importance of imputation is further demonstrated across various domains. Zemicheal and Dietterich (2019) showed that anomaly detection methods combined with data imputation techniques achieve significantly better performance than those without imputation. Similarly, Fan and Chow (2017c) demonstrated that clustering incomplete data can also benefit from missing data imputation. Furthermore, Tu et al. (2019) studied the problem of causal discovery in the presence of missing data, highlighting its relevance in causal inference.
@@
Certain problems can be naturally formulated as missing data imputation tasks. A primary example is found in recommendation systems, where an incomplete user-item interaction matrix (see Figure 6) is used; predicting the missing interactions directly yields user or item recommendations. Another example is image or video inpainting, which involves replacing corrupted or unwanted pixels or patches with plausible values. This technique is widely used in multimedia entertainment and film-making. A further instance is transductive classification (Goldberg et al., 2010), particularly when feature sets contain missing values. In this context, Goldberg et al. (2010) treated unknown labels as missing data and employed low-rank matrix completion to simultaneously predict the labels and impute the missing features.

### v2

@@ 1 Introduction
Missing data is a prevalent challenge across numerous fields, including social science, e-commerce, healthcare, bioinformatics, and communications. For instance, in social science surveys, respondents may refuse to answer certain questions, leading to non-responses (little1989analysis). In bioinformatics, gene expression matrices often contain significant missing values, or ‘dropouts’, which occur when expressed transcripts are not fully detected and are consequently recorded as zeros (troyanskaya2001missing; huang2018saver). Similarly, in sensor networks, node failures can result in incomplete data collection (faizin2019review). In e-commerce, the user-item interaction matrix is typically highly sparse, and predicting its missing values forms the basis of recommendation systems (koren09). Even in computer vision, tasks such as watermark and occlusion removal can be framed as missing data imputation problems. Beyond imputation, two other primary approaches for handling missing data are deletion and ignorance. The deletion method involves discarding any data instances or features that contain missing values, proceeding with analysis only on the remaining complete dataset. A common implementation is to remove rows or columns with missing entries from a data matrix. A significant drawback of this approach is that it can drastically reduce the dataset size, potentially compromising the reliability of subsequent analysis. Moreover, it fails to leverage the partial information available in the incomplete observations. The ignorance approach, conversely, overlooks the presence of missing values during analysis, often by replacing them with a simple placeholder like zero. This method presupposes that the analytical or learning algorithms are robust to such missingness. Intuitively, this strategy is only viable when the proportion of missing data is sufficiently small.
@@
Research on handling missing data dates back to the 1930s (wilks1932moments; yates1933analysis). For instance, wilks1932moments investigated the estimation of means, variances, and covariances for a bivariate normal population from incomplete samples, comparing two methods: maximum likelihood estimation and an approach based on systems of independent estimates. This foundational work was later expanded by (lord1955estimation) and (10.2307/2280845), who addressed the more general problem of estimating parameters for multivariate normal populations with incomplete data. A key literature review on this topic in multivariate statistics was subsequently compiled by (afifi1966missing). A significant theoretical advance was made by (rubin1976inference), who established that directly ignoring the missing data process is inappropriate for likelihood-based inference unless the data are Missing at Random. Around the same time, dempster1977maximum introduced the Expectation-Maximization (EM) algorithm for obtaining maximum likelihood estimates from incomplete data, which has since become one of the most influential methods in the field. Parallel to these methodological developments, researchers such as kalton1982imputing, little1988missing, and little1989analysis focused on missing data problems in the social sciences. Among them, little1989analysis provided a detailed analysis of different missingness patterns, including general and special cases like univariate and monotone missingness. From the perspective of psychological research, roth1994missing later reviewed classical methods for missing data imputation.
@@
Since real-world data are often represented in matrix form, several scholars have proposed performing missing data imputation by leveraging the underlying structure of the matrix, a problem known as matrix completion (CandesRecht2009; mazumder2010spectral). A particularly important and useful structure is low-rankness, which arises from the fact that real data matrices often lie in a low-dimensional subspace (udell2019big). This low-rank prior enables the effective recovery of missing values. While numerous approaches to matrix completion exist, some researchers have extended the low-rank assumption to more general structures, such as unions of subspaces or manifolds, which can lead to high-rank or even full-rank matrices (ErikssonBalzanoNowak2011; FANNLMC; pmlr-v70-ongie17a; fan2020polynomial). For instance, fan2020polynomial assumed that the data are drawn from a union of polynomials. Matrix completion methods have found applications in data pre-processing, image inpainting, collaborative filtering (and recommendation systems more broadly) (koren09), and link prediction (menon2011link).
@@
Higher-order tensors, as natural extensions of matrices, are also ubiquitous in science and engineering. For example, a color image with three channels can be represented as a third-order tensor, and fMRI scans across multiple time points can form a fourth-order tensor. To impute missing values in such data, tensor completion has been extensively studied as an extension of matrix completion (gandy2011tensor; acar2011scalable; liu2012tensor; qin2022low). Tensor completion has applications in recommendation systems, image and video inpainting, and knowledge graph completion, among others. Compared to matrix completion, tensor completion typically involves higher time and space complexity, and the landscape of tensor decomposition models is more diverse. However, both theoretical and empirical studies have shown that tensor completion often significantly outperforms matrix completion applied to matricized tensors in terms of recovery accuracy.
@@
Recently, deep learning (lecun2015deep; goodfellow2016deep), a prominent branch of machine learning, has been increasingly applied to missing data imputation. Deep learning typically refers to deep artificial neural networks, which are capable of approximating highly complex functions, as established by the universal approximation theorems (pinkus1999approximation; sonoda2017neural; lu2017expressive). This capacity enables deep learning to capture the intricate structures inherent in real-world data, leading to impressive performance in imputation tasks. Perhaps the most well-known deep learning model for this purpose is the autoencoder (hinton2006reducing). In this architecture, the input typically consists of incomplete data, and the output is the reconstructed, complete data (FAN2017540; gondara2018mida; 8891716). Beyond autoencoders, the widely recognized generative adversarial network (GAN) (goodfellow2014generative) has also been adapted for imputation (yoon2018gain; li2019misgan; yoon2020gamin). For instance, GAIN, introduced by yoon2018gain, incorporates additional ‘hint’ information into the discriminator to enable its operation on incomplete data. These hints guide the generator to produce imputations that adhere to the true underlying data distribution. Diffusion models (sohl2015deep; ho2020denoising; yang2023diffusion) have also been leveraged for missing data imputation (tashiro2021csdi). More recently, the potential of Large Language Models (LLMs) for missing data imputation has been investigated. More details will be provided in Section 5.7.
@@ 2 Related Work
The period from 2000 to 2010 yielded several influential surveys on missing data. Specifically, pigott2001review summarized common methods, including complete-case analysis, available-case analysis, single imputation, and model-based methods for multivariate normal data, as well as maximum likelihood via the EM algorithm. schafer2002missing provided a detailed discussion of the Missing at Random assumption and evaluated the effectiveness of maximum likelihood and Bayesian multiple imputation. During this time, little2002statistical also published a seminal book on the statistical analysis of missing data, the most recent edition of which was released in 2019. Meanwhile, donders2006gentle offered a concise review that compared single and multiple imputation through numerical experiments. graham2009missing presented a practical overview of the missing data literature, covering theoretical foundations and methods like normal-model multiple imputation and maximum likelihood, while also addressing practical issues such as using auxiliary variables to enhance power and reduce bias. The subsequent decade saw a continuation of this review effort. In 2010, Enders C.K. authored an applied book on handling missing data, with a latest edition published in 2022 (enders2022applied). That same year, baraldi2010introduction explored the theoretical foundations of missing data analysis, summarized conventional techniques, and provided accessible explanations of maximum likelihood and multiple imputation. After 2010, there are also a few review papers on missing data (nakai2011review; eekhout2012missing; silva2014brief; shen2015missing; pratama2016review; gabrio2017handling; wu2019imputation). For instance, pratama2016review briefly reviewed a few methods for handling missing data in time series.
A number of recent reviews have explored imputation methods across various fields. zhang2022handling surveyed methods for environmental monitoring, while sethia2023review provided a brief overview that categorizes techniques into simple and multiple imputation. In a 2024 review of 46 studies on electronic health records, ren2024moving concluded that while machine learning methods show significant promise, no single approach offers a universally generalizable solution. More recently, chourib2025missing presented a short review that classified strategies into three principal categories: preprocessing techniques, graph-based imputation, and algorithms inherently tolerant to missing values, accompanied by a numerical evaluation of five methods. It is worth noting that numerous other publications are titled as reviews or surveys on missing data, but most of them have a very narrow focus (e.g., (moorthy2014review; faizin2019review; wu2019imputation; zhang2025missing)) and will not be discussed here.
@@
(roth1994missing) psychology
(pigott2001review) statistics
(troyanskaya2001missing) bioinformatics
(schafer2002missing) statistics
(little2002statistical) statistics
(graham2009missing) psychology
(porter2012missing) social science
(cheema2014review) education
(silva2014brief) data analytics
(shen2015missing) remote sensing
(pratama2016review) time series
(du2020missing) monitoring systems
(emmanuel2021survey) machine learning
(thomas2021systematic) machine learning
(joel2022review) machine learning
(adhikari2022comprehensive) internet of things
(nijman2022missing) medical
(liu2022handling) healthcare & machine learning
(sun2023deep) machine learning
(ren2024moving) healthcare & machine learning
(alwateer2024missing) machine learning
(zhang2024comprehensive) transportation
(little2024missing) psychology
(LE2025100720) healthcare
(benhamza2025comprehensive) healthcare
(chourib2025missing) healthcare
@@
Notation Throughout this paper, we use $x$ or $X$ to denote a scalar, use $\mathbf{x}$ and $\mathbf{X}$ to denote vector and matrix respectively, use $\mathcal{X}$ to denote a set, and use $\bm{\mathcal{X}}$ to denote a tensor. Moreover, $\mathbf{x}$ may also denote a general data sample, corresponding to an image, a text, or even a multivariate time series. An incomplete $\mathbf{x}$ is denoted as $\tilde{\mathbf{x}}$, while the imputed $\tilde{\mathbf{x}}$ is denoted as $\hat{\mathbf{x}}$. This convention also applies to $\mathbf{X}$ and $\bm{\mathcal{X}}$.
@@ 3.1 Mechanisms of Missingness
When data are classified as missing at random (MAR), the occurrence of missing data is systematically linked to the observed data, but not to the unobserved data (rubin1976inference; heitjan1996distinguishing; little2019statistical). More formally, following the previous definition of $\mathbf{X}$ and $\mathbf{M}$, we split $\mathbf{X}$ into two parts, i.e., $\mathbf{X}=(\mathbf{X}_{\mathrm{obs}},\mathbf{X}_{\mathrm{mis}})$, where $\mathbf{X}_{\mathrm{obs}}$ and $\mathbf{X}_{\mathrm{mis}}$ denote the observed values and missing values respectively. The data are MAR if $P(\mathbf{M}\mid\mathbf{X})=P(\mathbf{M}\mid\mathbf{X}_{\mathrm{obs}})$. For instance, in a survey, if men are more likely to withhold their weight than women, but this non-disclosure is not connected to the actual weights, then such data would be considered MAR. In other words, the likelihood of survey completion is associated with their gender (which is completely observed) but not their weight. MAR is a weaker assumption than MCAR and is more common in real-world data. However, handling MAR data is more complex and usually requires more sophisticated statistical techniques to avoid bias, such as multiple imputation or maximum likelihood estimation.
@@
Missing not at random (MNAR), also referred to as non-ignorable nonresponse, represents a type of data that is neither MAR nor MCAR. In instances where data is MNAR, the absence of data is systematically connected to the unseen data, meaning that the missingness is linked to events or factors that the researcher has not measured. To elaborate on the previous example, a weight registry might face MNAR data if participants who are overweight are more inclined to decline completing a weight survey. Another instance could be a survey including sensitive queries like personal income. If individuals with higher earnings are more prone to withhold their income information (resulting in missing data), the missingness of the income data correlates with the actual missing income values. This situation exemplifies MNAR. More details and subtypes of MNAR can be found in (little2019statistical; pereira2024imputation). Addressing MNAR is a challenging task and may yield skewed results if not appropriately managed.
@@ 3.2 Single Imputation and Multiple Imputation
In single imputation, each missing value is replaced with a single estimated value. The methods used to estimate the missing value can vary, including mean imputation, median imputation, regression imputation, and many others that will be discussed in the subsequent sections of this paper. While single imputation is simple and easy to implement, it has significant limitations. It doesn’t reflect the uncertainty about the imputations (since each missing value is replaced by a single value), which can lead to underestimated standard errors and overly confident statistical inferences (rubin1987multiple).
Multiple imputation addresses the limitations of single imputation by replacing each missing value with a set of plausible values, creating multiple complete datasets. These datasets are then analyzed separately, and the results are combined to produce estimates and confidence intervals that incorporate missing-data uncertainty. This process reflects the uncertainty about the right model to impute missing data and the randomness inherent in the data itself. As a result, multiple imputation provides more accurate and statistically valid results than single imputation (rubin2018multiple; rubin1996multiple; murray2018multiple). The book (rubin2018multiple) provided a very detailed introduction, application, and discussion of multiple imputation. Compared to single imputation, the major limitation of multiple imputation is the higher computation cost, especially when advanced imputation algorithms are used.
@@ 3.3 Goals of Missing Data Imputation
Missing data imputation not only plays an important role in data pre-processing, but also is the final goal of many real scenarios such as recommendation systems. Moreover, missing data imputation can be used to improve efficiency and reduce the cost of data acquisition and analysis.
@@
In many data analysis tasks—such as classification, regression, clustering, novelty detection, and causal inference—algorithms typically require complete data. The performance of these algorithms is often directly influenced by the quality of the imputed values. For example, a support vector machine trained on data where missing values are zero-filled might achieve 60% classification accuracy on a test set; this accuracy can improve to 80% when a reliable imputation algorithm is used. The importance of imputation is further demonstrated across various domains. dietterich2018anomaly showed that anomaly detection methods combined with data imputation techniques achieve significantly better performance than those without imputation. Similarly, fan2017sparse demonstrated that clustering incomplete data can also benefit from missing data imputation. Furthermore, pmlr-v89-tu19a studied the problem of causal discovery in the presence of missing data, highlighting its relevance in causal inference.
@@
Certain problems can be naturally formulated as missing data imputation tasks. A primary example is found in recommendation systems, where an incomplete user-item interaction matrix (see Figure 6) is used; predicting the missing interactions directly yields user or item recommendations. Another example is image or video inpainting, which involves replacing corrupted or unwanted pixels or patches with plausible values. This technique is widely used in multimedia entertainment and film-making. A further instance is transductive classification (NIPS2010_3932), particularly when feature sets contain missing values. In this context, NIPS2010_3932 treated unknown labels as missing data and employed low-rank matrix completion to simultaneously predict the labels and impute the missing features.